An ordinary deck of 52 cards of four suits. The queen of spades is randomly drawn and removed from the well shuffled deck. What is the conditional probability p that one card drawn randomly from the remaining deck will be a face card or a club?
Answer #1

Number of suits =4

Number of face cards(Jack, Queen, king ) in each suit = 3

Number of face cards in standard deck = 4 x3=12

A: Event of drawing queen of spades (first card)

B : Event of drawing a face card or a club(second card)

The queen of spades is randomly drawn and removed from the well shuffled deck.

p : Conditional probability hat one card drawn randomly from the remaining deck will be a face card or a club given that the queen of spades is randomly drawn and removed from the well shuffled deck = P(B|A)

Given that first card drawn is queen of spades and removed

Queen of spades is face card

Number of cards left in the deck = 52-1 =51

Number of face cards left in the deck = 12-1 = 11

Number of club cards left in the deck = 13

Number of club and face cards left in the deck = 3 (Club jack, Club queen, club king)

Probability of drawing a face card or club card = Probability of drawing a face card + Probability of drawing a club card - Probability of drawing club face card

Probability of drawing a face card = Number of face cards left in the deck / Number of cards left in the deck =11/51

Probability of drawing a club card = Number of club cards left in the deck / Number of cards left in the deck =13/51

Probability of drawing a club card - Probability of drawing club face card = Number of club and face cards left in the deck /  Number of cards left in the deck = 3/51

Probability of drawing a face card or club card = Probability of drawing a face card + Probability of drawing a club card - Probability of drawing club face card = 11/51 + 13/ 51 - 3/51 = 21/51

p : Conditional probability hat one card drawn randomly from the remaining deck will be a face card or a club given that the queen of spades is randomly drawn and removed from the well shuffled deck = 21/51 = 7/17

p : Conditional probability hat one card drawn randomly from the remaining deck will be a face card or a club given that the queen of spades is randomly drawn and removed from the well shuffled deck = 7/17 = 0.411765